Determine the solution set of (x + 2)2 = 36.
{6, -8}
{4, -8}
{4, -6}

Answer by colliefan(242)

What you need to remember is that a positive number and it's additive inverse both give you the same number when they are squared. Eg; 4 and -4 both give you 16 when they are squared.
The square root of 36 is 6. The solutions for x then must be either 6 squared or -6 squared.
x+2=6 and x+2=-6 then are equations that give the two solutions.
x+2=6
x=4
x+2=-6
x=-8
(4,-8)